OpenStudy (anonymous):

how do u get the area of 125,788 ft to be 10% and 20% larger?

OpenStudy (anonymous):

multiply by \[1.1\] and \[1.2\] respectively. is this the entire problem?

OpenStudy (anonymous):

can i ask why its 1.1 and 1.2 but not just 0.1 and 0.2?

OpenStudy (anonymous):

if you multiply by .1 or .2 you will find 10% or 20%, but you want to INCREASE it by 10%

OpenStudy (anonymous):

so if you like, you could find 10% and then add it on, or else you can just multiply by 1.1 here we can try it

OpenStudy (anonymous):

ohhhhh! i feel quite stupid at the moment.......but thank u! so its 138,366.8 and 150,945.6?

OpenStudy (anonymous):

\[125,788(1.1)= 138366.8\] or \[125,788+.1\times 125,788=125,788+12,578.8=138366.8\]

OpenStudy (anonymous):

OMG! thank u sooo much!

OpenStudy (anonymous):

works out the same either way, yw

OpenStudy (anonymous):

why did you become my fan? lols

OpenStudy (anonymous):

yup! :) is that okay?

OpenStudy (anonymous):

i asked why. lol and its fine. i dont care. haha

OpenStudy (anonymous):

alrighty!